# Trainable Dynamic Mask Sparse Attention

### v1

@@ 6.1 Limitations of Existing Approaches
Post-hoc Sparsification Degradation.
The performance degradation caused by post-hoc sparsification stems from the fundamental mismatch between existing methods and the optimization trajectory of pretrained models. As demonstrated by Chen et al. [8], retaining only the top 20% of attention weights covers only 70% of the total attention scores. This forced sparsification strategy compels models to deviate from the optimal parameter configurations learned on large-scale corpora. More critically, this approach causes irreversible damage to key structural components in pretrained models, such as retrieval heads and copy heads, as these specialized attention heads are misidentified as "unimportant" and pruned during inference.
Training-Inference Efficiency Gap.
Most sparse attention methods optimize only for inference, neglecting training-phase computational demands. This creates bottlenecks across LLM development: pretraining on long documents, long-context fine-tuning, and reinforcement learning. Without effective training-time sparsity support, these crucial phases remain constrained by $O(n^{2})$ computational complexity, limiting development of more capable long-context models.
Non-differentiable Components and Inefficient Backpropagation.
Non-differentiable components and inefficient backpropagation problems reveal the technical shortcomings of existing methods in terms of trainability. The discrete operations in methods like ClusterKV [33] and MagicPIG [8] introduce discontinuities in computational graphs, which block gradient flow and hinder the learning of optimal sparse patterns. Even trainable methods like HashAttention [15] suffer from memory access inefficiencies due to token-granular selection, which is incompatible with the contiguous memory access and block-wise computation requirements of efficient attention techniques, such as FlashAttention. Consequently, these implementations are forced to revert to naive implementations with low hardware utilization, significantly degrading training efficiency.
@@ 6.2 How Dynamic Mask Attention Addresses Core Issues
Native trainable sparsity is DMA’s key innovation for addressing post-hoc sparsification issues. Unlike traditional methods, DMA embeds sparsity into the model architecture from the ground up, ensuring that sparse attention patterns are fully aligned with the model’s optimization trajectory. Specifically, DMA retains complete, uncompressed KV caches $k=\operatorname{concat}([k_{1},\ldots,k_{t}])$ and $v=\operatorname{concat}([v_{1},\ldots,v_{t}])$, ensuring the original fidelity of historical information and precise recall capabilities, avoiding information bottlenecks that may arise from fixed-state compression in State Space Models. This comprehensive information retention mechanism enables DMA to precisely access any token in the historical context at any moment, without losing critical information due to lossy compression methods like Mamba. More importantly, DMA’s sparsification occurs during the attention weight computation phase, rather than in post-training processing, ensuring that models do not deviate from pre-trained parameter configurations during sparsification, thereby protecting key structural components, such as retrieval heads and copy heads, from damage.
Unified Training-Inference Architecture.
@@
The unified training-inference architecture eliminates the fundamental gap in training-inference efficiency that exists in existing methods. DMA’s dynamic weight computation $\delta=\exp(\tau(v\Delta)\times A)$ uses identical sparsification strategies during both training and inference phases. This consistency ensures that models can learn optimal sparse patterns during training and seamlessly apply these patterns during inference. This unified architecture particularly benefits three critical stages of modern LLM development: the pretraining stage can efficiently process long document sequences; the long-context fine-tuning stage can adapt to specific task requirements; the reinforcement learning stage can effectively update attention weights through policy gradients. DMA reduces computational complexity from $O(n^{2})$ to $O(n\cdot w)$, enabling the training of larger-scale long-context models.
Fully Differentiable Design.
@@
The fully differentiable design ensures that DMA maintains gradient flow continuity throughout the entire computation process. The computation of dynamic mask weights $\delta$ is based entirely on differentiable operations, including linear transformations of value representations, non-negative activation functions $\tau(\cdot)$, and exponential functions, thereby avoiding gradient interruptions caused by discrete operations such as k-means clustering and SimHash. Although the mask generation process involves topk operations, since it is not the core learning objective of DMA but merely a tool for sparse selection, we can maintain practical gradient propagation through continuous relaxation techniques or straight-through estimators. Moreover, the attention weight computation part is designed such that the gradients for masked positions should naturally be zero, so skipping computation and setting gradients to zero is the correct behavior. This design enables the model to learn optimal attention patterns that are sparse in an end-to-end manner, dynamically adjusting which historical positions are most critical for current reasoning, thereby achieving truly content-aware, selective computation. Additionally, each head in a multi-head attention mechanism can independently generate different sparse patterns, thereby maximizing the representational capabilities of the multi-head architecture by focusing on different information segments in distinct subspaces.
@@ 6.3 Limitations and Future Works
Adaptive window size selection is the primary challenge facing DMA. While the current fixed window size design provides predictable computational complexity, it may not optimally adapt to the dynamic demands of different tasks and contexts. For instance, code generation tasks may require larger windows to capture long-range structural dependencies, while simple question-answering tasks may only need smaller windows. Future research directions include developing adaptive window size selection mechanisms based on task complexity, sequence length, and content features, potentially through reinforcement learning or meta-learning approaches to dynamically optimize window parameters. Alternatively, designing hierarchical multi-scale attention structures can be considered to capture dependencies across different ranges simultaneously.
Position Encoding Enhancement.
@@
Our needle-in-a-haystack experiments revealed an intriguing phenomenon: trainable sparse attention mechanisms, such as DMA, exhibit stronger length extrapolation capabilities compared to dense attention when context lengths exceed the pretraining bounds. This finding suggests that the fundamental bottleneck for extrapolation may lie in the position encoding method rather than the attention mechanism itself. Current RoPE-based position encodings struggle with out-of-distribution sequence lengths, but DMA’s dynamic sampling architecture offers a potential alternative pathway for encoding positional information. Specifically, the zero-order hold sampling values that are added as attention biases can be explored to explicitly incorporate positional information into these sampling values, potentially replacing or complementing RoPE to create a more extrapolation-friendly encoding scheme. Such an approach might leverage the inherent advantages of sparse attention’s selective computation to create position representations that scale more naturally to unseen lengths. This direction could help address one of the most persistent challenges in long-context modeling: maintaining consistent positional understanding across arbitrary sequence lengths without requiring length-specific fine-tuning.
Multi-Modal Extension.
@@
Multi-modal extension represents an essential direction for DMA development. The current DMA design is primarily optimized for text sequences; however, modern AI systems increasingly require processing mixed inputs of text, images, audio, and video. Attention sparsity in multi-modal scenarios exhibits more complex patterns: interactions between different modalities may require different attention distributions, temporally aligned multi-modal information may need synchronized attention mechanisms, and modality-specific long-range dependencies may require specialized sparse patterns. Future research can explore modality-aware dynamic mask generation, coordination mechanisms for cross-modal attention weights, and specialized sparse pattern designs for different modal characteristics.
@@ Acknowledgments
We would like to express our gratitude to the OpenSeek project team at Beijing Academy of Artificial Intelligence for their support in developing the hardware kernels. We would also like to thank Professor Albert Gu from Carnegie Mellon University for his valuable guidance and suggestions in connecting the concepts of state-space and self-attention.
@@ Appendix A Dynamic Mask Attention Implementation
Figure 8: Dynamic Mask Attention implementation in PyTorch
@@ Appendix B Experiment Setup
Algos Params Steps Batch LR $n_{layers}$ $d_{model}$ $n_{h}$ $w$ $d_{c}$ $B$ $B^{\prime}$ $k$
MHA $\approx$ 80M 13,500 0.128M tokens 3e-3 12 768 6 - - - - -
SWA $\approx$ 80M 13,500 0.128M tokens 3e-3 12 768 6 1024 - - - -
MLA $\approx$ 80M 13,500 0.128M tokens 3e-3 12 768 6 1024 192 - - -
NSA $\approx$ 80M 13,500 0.128M tokens 3e-3 12 768 6 512 192 32 64 16
DMA $\approx$ 80M 13,500 0.128M tokens 3e-3 12 768 6 1024 - - - -
MHA $\approx$ 200M 20,800 0.192M tokens 2e-3 16 1024 8 - - - - -
SWA $\approx$ 200M 20,800 0.192M tokens 2e-3 16 1024 8 1024 - - - -
MLA $\approx$ 200M 20,800 0.192M tokens 2e-3 16 1024 8 - 256 - - -
NSA $\approx$ 200M 20,800 0.192M tokens 2e-3 16 1024 8 512 192 32 64 16
DMA $\approx$ 200M 20,800 0.192M tokens 2e-3 16 1024 8 1024 - - - -
MHA $\approx$ 680M 35,000 0.392M tokens 1e-3 24 1536 12 - - - - -
NSA $\approx$ 680M 35,000 0.392M tokens 1e-3 24 1536 12 512 192 32 64 16
DMA $\approx$ 680M 35,000 0.392M tokens 1e-3 24 1536 12 1024 - - - -
MHA $\approx$ 1.7B 40,000 1M tokens 1e-3 28 2048 16 - - - - -
NSA $\approx$ 1.7B 40,000 1M tokens 1e-3 28 2048 16 512 256 32 64 16
DMA $\approx$ 1.7B 40,000 1M tokens 1e-3 28 2048 16 2048 - - - -
Appendix C Dynamic Mask Attention Hatmaps
Figure 9: Dynamic Mask Attention Heatmaps. The heatmaps show the attention weights of each head in the Dynamic Mask Attention mechanism, indicating which tokens each head focuses on.

### v6

@@ 7.1 Limitations of Existing Approaches
Post-hoc Sparsification Degradation. The performance degradation caused by post-hoc sparsification stems from the fundamental mismatch between existing methods and the optimization trajectory of pretrained models. As demonstrated by Chen et al. [chen2024magicpig], retaining only the top 20% of attention weights covers only 70% of the total attention scores. This forced sparsification strategy compels models to deviate from the optimal parameter configurations learned on large-scale corpora. More critically, this approach causes irreversible damage to key structural components in pretrained models, such as retrieval heads and copy heads, as these specialized attention heads are misidentified as "unimportant" and pruned during inference. This structural destruction directly leads to significant performance degradation in tasks that require precise information retrieval and copying.
Training-Inference Efficiency Gap. Most sparse attention methods optimize only for inference, neglecting training-phase computational demands. This creates bottlenecks across LLM development: pretraining on long documents, long-context fine-tuning, and reinforcement learning. Without effective training-time sparsity support, these crucial phases remain constrained by $O(n^{2})$ computational complexity, limiting development of more capable long-context models.
Non-differentiable Components and Inefficient Backpropagation. Non-differentiable components and inefficient backpropagation problems reveal the technical shortcomings of existing methods in terms of trainability. The discrete operations in methods like ClusterKV [liu2024clusterkv] and MagicPIG [chen2024magicpig] introduce discontinuities in computational graphs, which block gradient flow and hinder the learning of optimal sparse patterns. Even trainable methods like HashAttention [desai2024hashattention] suffer from memory access inefficiencies due to token-granular selection, which is incompatible with the contiguous memory access and block-wise computation requirements of efficient attention techniques, such as FlashAttention. Consequently, these implementations are forced to revert to naive implementations with low hardware utilization, significantly degrading training efficiency.
@@ 7.2 How Dynamic Mask Attention Addresses Core Issues
Native Trainable Sparsity. Native trainable sparsity is DMA’s key innovation for addressing post-hoc sparsification issues. Unlike traditional methods, DMA embeds sparsity into the model architecture from the ground up, ensuring that sparse attention patterns are fully aligned with the model’s optimization trajectory. Specifically, DMA retains complete, uncompressed KV caches $k=\operatorname{concat}([k_{1},\ldots,k_{t}])$ and $v=\operatorname{concat}([v_{1},\ldots,v_{t}])$, ensuring the original fidelity of historical information and precise recall capabilities, avoiding information bottlenecks that may arise from fixed-state compression in State Space Models. This comprehensive information retention mechanism enables DMA to precisely access any token in the historical context at any moment, without losing critical information due to lossy compression methods like Mamba. More importantly, DMA’s sparsification occurs during the attention weight computation phase, rather than in post-training processing, ensuring that models do not deviate from pre-trained parameter configurations during sparsification, thereby protecting key structural components, such as retrieval heads and copy heads, from damage.
@@
Unified Training-Inference Architecture. The unified training-inference architecture eliminates the fundamental gap in training-inference efficiency that exists in existing methods. DMA’s dynamic weight computation $\delta=\exp(\tau(v\Delta)\times A)$ uses identical sparsification strategies during both training and inference phases. This consistency ensures that models can learn optimal sparse patterns during training and seamlessly apply these patterns during inference. This unified architecture particularly benefits three critical stages of modern LLM development: the pretraining stage can efficiently process long document sequences; the long-context fine-tuning stage can adapt to specific task requirements; the reinforcement learning stage can effectively update attention weights through policy gradients. DMA reduces computational complexity from $O(n^{2})$ to $O(n\cdot w)$, enabling the training of larger-scale long-context models.
@@
Fully Differentiable Design. The fully differentiable design ensures that DMA maintains gradient flow continuity throughout the entire computation process. The computation of dynamic mask weights $\delta$ is based entirely on differentiable operations, including linear transformations of value representations, non-negative activation functions $\tau(\cdot)$, and exponential functions, thereby avoiding gradient interruptions caused by discrete operations such as k-means clustering and SimHash. Although the mask generation process involves a top-k operation, it is not the core learning objective of DMA but merely a tool for sparse selection; thus, we only use this discrete operation in the forward pass. Moreover, the attention weight computation part is designed such that the gradients for masked positions should naturally be zero, so skipping computation and setting gradients to zero is the correct behavior. This design enables the model to learn optimal attention patterns that are sparse in an end-to-end manner, dynamically adjusting which historical positions are most critical for current reasoning, thereby achieving truly content-aware, selective computation. Additionally, each head in a multi-head attention mechanism can independently generate different sparse patterns, thereby the representational capabilities of the multi-head architecture by focusing on different information segments in distinct subspaces.
@@ 7.3 Limitations and Future Works
Adaptive Window Size Selection. Adaptive window size selection is the primary challenge facing DMA. While the current fixed window size design provides predictable computational complexity, it may not optimally adapt to the dynamic demands of different tasks and contexts. For instance, code generation tasks may require larger windows to capture long-range structural dependencies, while simple question-answering tasks may only need smaller windows. Future research directions include developing adaptive window size selection mechanisms based on task complexity, sequence length, and content features, potentially through reinforcement learning or meta-learning approaches to dynamically optimize window parameters. Alternatively, designing hierarchical multi-scale attention structures can be considered to capture dependencies across different ranges simultaneously.
@@
Position Encoding Enhancement. Our needle-in-a-haystack experiments revealed an intriguing phenomenon: trainable sparse attention mechanisms, such as DMA, exhibit stronger length extrapolation capabilities compared to dense attention when context lengths exceed the pretraining bounds. This finding suggests that the fundamental bottleneck for extrapolation may lie in the position encoding method rather than the attention mechanism itself. Current RoPE-based position encodings struggle with out-of-distribution sequence lengths, but DMA’s dynamic sampling architecture offers a potential alternative pathway for encoding positional information. Specifically, the zero-order hold sampling values that are added as attention biases can be explored to explicitly incorporate positional information into these sampling values, potentially replacing or complementing RoPE to create a more extrapolation-friendly encoding scheme. Such an approach might leverage the inherent advantages of sparse attention’s selective computation to create position representations that scale more naturally to unseen lengths. This direction could help address one of the most persistent challenges in long-context modeling: maintaining consistent positional understanding across arbitrary sequence lengths without requiring length-specific fine-tuning.
@@
Multi-Modal Extension. Multi-modal extension represents an essential direction for DMA development. The current DMA design is primarily optimized for text sequences; however, modern AI systems increasingly require processing mixed inputs of text, images, audio, and video. Attention sparsity in multi-modal scenarios exhibits more complex patterns: interactions between different modalities may require different attention distributions, temporally aligned multi-modal information may need synchronized attention mechanisms, and modality-specific long-range dependencies may require specialized sparse patterns. Future research can explore modality-aware dynamic mask generation, coordination mechanisms for cross-modal attention weights, and specialized sparse pattern designs for different modal characteristics.
Integration with Modern Frameworks. Seamless integration of DMA into mainstream deep learning frameworks, such as PyTorch and the Hugging Face Transformers library, is crucial for its widespread adoption and practical impact. This requires developing a user-friendly and highly optimized implementation that can be easily incorporated into existing model architectures and training pipelines. A key aspect of this integration is the development of efficient, low-level kernels, potentially using Triton or CUDA, to ensure that the performance benefits of DMA are fully realized on modern hardware. Providing a plug-and-play module compatible with the Hugging Face Transformers library would significantly lower the barrier for researchers and practitioners to apply DMA to their models and tasks, thereby fostering further innovation and comparative studies in the field of sparse attention.
@@ Acknowledgments
We would like to express our gratitude to the OpenSeek project team at Beijing Academy of Artificial Intelligence for their support in developing the hardware kernels. We would also like to thank Professor Albert Gu from Carnegie Mellon University for his valuable guidance and suggestions in connecting the concepts of state-space and self-attention. Additionally, we extend our thanks to all the friends in the community who provided feedback and suggestions, your support has been instrumental in the continuous improvement of this work.
@@ Appendix A Dynamic Mask Attention Implementation
Listing 1: Dynamic Mask Attention implementation in PyTorch
@@ Appendix B Experiment Setup
To make the comparison between attention variants fair and reproducible, we standardize the model, data pipeline, optimization, and evaluation, only changing the attention module and its related hyperparameters. All experiments were conducted using the open-source PyTorch images [nv2022pytorch] and the Transformers framework [wolf-etal-2020-transformers]. We use SmolLMCorpus [benallal2024smollmcorpus] as training data. For evaluation frameworks, we utilized the LM evaluation harness [eval-harness] from EleutherAI for perplexity tasks, and the lighteval [Clémentine2023lighteval] from HuggingFace for downstream tasks. Table 4 lists the model sizes and key hyperparameters used at each scale. We keep the depth, width, and number of heads consistent across variants under the same parameter budget, only changing the specific parameters of the attention variants, with parameter symbols consistent with those in the original papers of the attention variants.
We summarize the meaning of the columns in Table 4 and clarify which hyperparameters are used by each attention variant.
  • Params: total number of model parameters.
  • Steps: training steps.
  • Batch: tokens per step.
  • LR: peak learning rate.
  • $n_{layers}$: number of Transformer layers.
  • $d_{model}$: model hidden size.
  • $n_{h}$: number of attention heads.
  • $n_{h_{kv}}$: number of KV heads. In all our configurations we set $n_{h_{kv}}=n_{h}/2$.
  • MHA (full attention): standard scaled dot-product attention. Sparse-specific columns ($w$, $d_{c}$, $B$, $B^{\prime}$, $k$) are not used.
  • SWA (sliding-window attention): $w$ is the sliding window size; other sparse-specific columns are not used.
  • MLA (multi-head latent attention): $d_{c}$ is the latent/compression dimension; other columns are not used.
  • NSA (native sparse attention): $d_{c}$ is the compression dimension, $w$ is the sliding window size, $B$ is the compressing block size, $B^{\prime}$ is the selection block size, and $k$ is the num selected blocks. All settings following the original paper.
  • DMA (dynamic mask attention): $w$ is the per-head top-$w$ kept keys for dynamic masks; other sparse-specific columns are not used.
@@
Algos Params Steps Batch LR $n_{layers}$ $d_{model}$ $n_{h}$ $n_{h_{kv}}$ $w$ $d_{c}$ $B$ $B^{\prime}$ $k$
MHA $\approx$ 80M 13,500 0.128M tokens 3e-3 12 768 6 3 - - - - -
SWA $\approx$ 80M 13,500 0.128M tokens 3e-3 12 768 6 3 1024 - - - -
MLA $\approx$ 80M 13,500 0.128M tokens 3e-3 12 768 6 3 - 192 - - -
NSA $\approx$ 80M 13,500 0.128M tokens 3e-3 12 768 6 3 512 192 32 64 16
DMA $\approx$ 80M 13,500 0.128M tokens 3e-3 12 768 6 3 1024 - - - -
MHA $\approx$ 200M 20,800 0.192M tokens 2e-3 16 1024 8 4 - - - - -
SWA $\approx$ 200M 20,800 0.192M tokens 2e-3 16 1024 8 4 1024 - - - -
MLA $\approx$ 200M 20,800 0.192M tokens 2e-3 16 1024 8 4 - 256 - - -
NSA $\approx$ 200M 20,800 0.192M tokens 2e-3 16 1024 8 4 512 192 32 64 16
DMA $\approx$ 200M 20,800 0.192M tokens 2e-3 16 1024 8 4 1024 - - - -
MHA $\approx$ 680M 35,000 0.392M tokens 1e-3 24 1536 12 6 - - - - -
NSA $\approx$ 680M 35,000 0.392M tokens 1e-3 24 1536 12 6 512 192 32 64 16
DMA $\approx$ 680M 35,000 0.392M tokens 1e-3 24 1536 12 6 1024 - - - -
MHA $\approx$ 1.7B 40,000 1M tokens 1e-3 28 2048 16 8 - - - - -
NSA $\approx$ 1.7B 40,000 1M tokens 1e-3 28 2048 16 8 512 256 32 64 16
DMA $\approx$ 1.7B 40,000 1M tokens 1e-3 28 2048 16 8 2048 - - - -
Table 5: Speed Benchmark Configurations. The common settings for running time curves and the sparse hyperparameters for each method.
Algo warmups runs $n_{h}$ $n_{h_{kv}}$ $d_{h}$ $w$ $d_{c}$ $B$ $B^{\prime}$ $k$ precision
MHA11footnotemark: 1 3 1,000 32 8 128 - - - - - bf16
SWA22footnotemark: 2 3 1,000 32 8 128 1024 - - - - bf16
MLA33footnotemark: 3 3 1,000 32 8 128 - 192 - - - bf16
NSA44footnotemark: 4 3 1,000 32 8 128 512 256 32 64 16 bf16
DMA 3 1,000 32 8 128 1024 - - - - bf16
Appendix C Attention Heatmaps
Figure 11: Full Heatmaps of Dynamic Mask Attention. The heatmaps show the attention weights of each head in the Dynamic Mask Attention mechanism, indicating which tokens each head focuses on.